OpenStudy (anonymous):

Find the product: 4 3/8 / 2 1/2

jhonyy9 (jhonyy9):

35/8/5/2 =35/8 * 2/5 = 7/4

OpenStudy (anonymous):

10 15/16

hero (hero):

It said "find product" That threw me off

jhonyy9 (jhonyy9):

Hero check the question please

jhonyy9 (jhonyy9):

how is there write

hero (hero):

jhonyy9 it's not me, it's the question that's in conflict. She has "find product" posted but she separates two fractions with a division sign. I don't believe she understands how confusing the question is. You shouldn't be too confident with your answer either.

jhonyy9 (jhonyy9):

yes there is write product but like sign there is divide

OpenStudy (anonymous):

She should type it in wolfram and show us how it looks from there.

jhonyy9 (jhonyy9):

ok but one math exercise we solve after how have there wrote and not after the text

jhonyy9 (jhonyy9):

Kyrie --- ATTENTION PLEASE !!!

OpenStudy (anonymous):

Oh! yeah? (:

OpenStudy (anonymous):

\[4 \left(\begin{matrix}3 \\ 8\end{matrix}\right) \div 2 \left(\begin{matrix}1 \\ 2\end{matrix}\right)=\]